Suppose an economy is in recession with a negative GDP gap of $800 billion. If MPC is 0.9, how much should the government spend to eliminate the negative GDP gap?
1
Expert's answer
2018-11-12T15:42:09-0500

m = 1/(1-МРС)=1/(1-0.9)=10

ΔDG =ΔGDP/m=800/10=$80 billion.

Answer: the government should spend $80 billion to eliminate the negative GDP gap.